morrocco has a population of 32 million adults, of which 24 million own cellular phones. If six Morroccans adults are randomly selected, what the probability that exactly three own a cellular phone?
1
Expert's answer
2019-05-24T05:57:30-0400

In accordance with the Bernoulli formula, the probability of k successes from n tries (with a probability of success p) is P(n;k;p)=Cknpk(1-p)n-k.

For this case: n=6, k=3; p=24000000/32000000=24/32=3/4.

P(6;3;0.75)=C360.753(1-0.75)6-3=6!/(3!(6-3)!)0.7530.253=6*5*4/6*0.7530.253=20*0.7530.253=0.131836